OpenStudy (anonymous):

Arrange these numbers from least to greatest: -5/6,..-4/9,..-3/5,..-4/7

OpenStudy (tkhunny):

On occasion, the decimal expression is the easiest... -5/6 = -0.83333... -4/9 = -0.44444... -3/5 = -0.60000... -4/7 = -0.57142... What say you? Any easier?